Question 3 Consider a closed economy described by the following equations: Y=C+I+G Y-5,000 G 1,000 T= 1,000 C 250+0.75 (Y -T) 1,000-50 a. (3 points) In this economy, compute private saving, public saving, and national saving. b. (2 points) Find the equilibrium interest rate. c. (2 points) Draw a graph containing the saving and investment curves for this economy Show the financial market equilibrium. d. (2 points) Now suppose the G rises to 1,250. Compute private saving, public saving, and national saving. (2 points) Find the new equilibrium interest rate. (1 points) On a new graph, draw the investment curve and the new saving curve. Show e. f. the new equilibrium in the financial market.
Answer #1

Y = C+I+G

Y = 5000

G = 1000

T = 1000

C = 250+0.75(Y-T) = 250+0.75(5000-1000)

C = 3250

I = 1000-50r

a) Private savings = Y --T -- C = 5000-1000-3250 = 750

Public savings = T-G = 1000-1000= 0

National savings = 750+0=750

b) Y = C+I+G

5000= 3250+1000-50r+1000

50r = 5250-5000

50r = 250

r = 5%

d) If G = 1250

Private savings = Y-T-C = 5000-1000-3250 = 750

Public savings = T -G = 1000-1250 = -250

National savings = 750-250 = 500

e) Y = C+I+G

5000 = 3250+1000-50r+1250

5000 = 5500-50r

50r=500

r = 10%
Answer #2
Formula of equilibrium interest rate ?
answered by: BCM20127 Amaan
Answer #3

SOLUTION 


Y = C + I + G 

C = 250 + 0.75(Y - T)

I = 1000 - 50 * r

=> Y = [ 250 + 0.75(Y - T) ] + [ 1000 - 50 * r ] + G

=> Y = 250 + 0.75Y - 0.75T + 1000 - 50 * r + G

=> (Y - 0.75Y) = 1250 - 0.75T - 50 * r + G 

=> 0.25 Y = 1250 - 0.75T - 50 * r + G

=> 0.25 * 5000 = 1250 - 0.75 * 1000 - 50 * r + 1000

=> 50 * r = 250 

=> r = 5 


So, 


C = 250 + 0.75(500 - 1000) = 3250


I = 1000 - 50 * 5 = 750 



a.


Private saving = Y - C - T = 5000 - 3250 - 1000 = 750 (ANSWER)


Public saving = T - G = 1000 - 1000 = 0


National saving = Private saving + Public saving = 750 = 0 = 750 (ANSWER).


b.


Equilibrium interest rate = 5 % (ANSWER). (As calculated above).


d.


G = 1250.


Private saving = Y - C - T = 5000 - 3250 - 1000 = 750 (ANSWER)


Public saving = T - G = 1000 - 1250 = - 250 (ANSWER)


National saving = 750 + (- 250) = 500 (ANSWER).


e.


I = Y - C - G = 5000 - 3250 - 1250 = 500 

I = 1000 - 50 * r

=> 500 = 1000 - 50 * r

=> r = 10 % (ANSWER).

answered by: Tulsiram Garg